OpenStudy (anonymous):

Suppose that a sledgehammer that weighs 18 pounds on Earth will weigh about 4 pounds on the planet Namar. Based on this, your weight on Namar would be about ___ times your weight on Earth. A. 0.25 B. 4 C. 0.5 D. Not enough information.

OpenStudy (johnweldon1993):

Well you know that clearly it is less than the weight on Earth...so it can't be more than 1....so 4 is automatically out of the question To find what the weight would be...you would do \[\frac{ 4 }{ 18 } = ?\]

OpenStudy (anonymous):

0.25?

OpenStudy (johnweldon1993):

Right..that is the closest to the true answer...so yes it would be correct

OpenStudy (anonymous):

Will you help me with the next one?

OpenStudy (anonymous):

A sledgehammer weighs 18 pounds on Earth and 4 pounds on the planet Namar. Based on the numbers, an 80–pound earthling would weigh about _____ pounds on Namar. 4 320 18 40

OpenStudy (johnweldon1993):

For this one we can use a proportion... \[\frac{ 18 }{ 4 } = \frac{ 80 }{ x }\] We would cross multiply and solve for x

OpenStudy (anonymous):

so 18?

OpenStudy (anonymous):

Just one more sorry if im being a bother. i just need to pass this class. An astronaut weighs 170 pounds on Earth and 400 pounds on Jupiter. The astronaut captures an alien that weighs 120 pounds on Jupiter. What is the alien's weight on Earth? 34 pounds 51 pounds 78 pounds 283 pounds

OpenStudy (johnweldon1993):

Yes 18 good.. and no problem dont worry about it....and okay again this can be a proportion \[\frac{ 170 }{ 400 } = \frac{ x }{ 120 }\] Cross multiply and solve for x again

OpenStudy (anonymous):

51?

OpenStudy (johnweldon1993):

Correct again!

OpenStudy (anonymous):

okay so i think i have the answer for this one but to sdouble check An astronaut weighs 170 pounds on Earth but only 68 pounds on Mars. While visiting the planet, the astronaut removes a rock sample that weighs 25 pounds on Mars. What is the weight of the rock sample on Earth? 1.5 pounds 6.25 pounds 10 pounds 62.5 pounds

OpenStudy (anonymous):

i got 10 pounds. but i feel like it would be wrong.

OpenStudy (johnweldon1993):

Yeah that is wrong...show me what you did for it...

OpenStudy (anonymous):

i used proportion. \[\frac{ 170 }{ 60 }=\frac{x }{ 25 }\]

OpenStudy (anonymous):

i set the proportion up wrong didnt i?

OpenStudy (johnweldon1993):

No you set it up correctly...but used the wrong number \[\frac{ 170 }{ 68 } = \frac{ x }{ 25 }\] cross multiply and solve for 'x'

OpenStudy (anonymous):

62.5?

OpenStudy (johnweldon1993):

Right there you go

OpenStudy (anonymous):

thank you so much!!

OpenStudy (johnweldon1993):

No problem! Good work!